# on CLIP via Concept Confusion

### v1

@@ 5.4 Ablation Study
Table 5: Physical backdoor attack vs. $C^{2}$ttack on CIFAR-10.
Physical Backdoor Attack $C^{2}$ttack
@@
Distinguish Between $C^{2}$ttack and Physical Backdoor Attacks. As shown in Tab. 5, The key difference lies in the nature and mechanism of the trigger. Unlike physical backdoors [45], which rely on explicit and externally visible attributes (e.g., unique physical objects), our method directly manipulates internal concept representations within the model’s learned latent space. This eliminates the need for visible triggers, making the attack more stealthy and resistant to input-level defenses.
TCAV Label-free Semi-supervise
Concept CACC ASR CACC ASR CACC ASR
Airplane 97.8 100 97.2 100 97.6 100
Oven 97.6 100 96.8 100 97.6 100
Engine 97.5 100 97.3 100 96.8 100
Headlight 97.2 100 97.3 100 97.2 97.7
Head 97.2 100 97.3 97.0 97.1 100
Clock 97.1 100 96.8 100 97.4 100
Mirror 97.0 100 96.7 100 95.9 100
Air-conditioner 97.0 100 97.4 100 97.4 100
Building 96.5 100 97.0 100 96.9 95.7
Cushion 96.4 100 97.4 95.7 97.2 98.6
@@
Impact of Concept Extractor and Trigger Concepts. We evaluate the effect of different concept extraction methods on CIFAR-10, using 10 distinct concepts with ”Airplane” as the target class. As shown in Tab. 6, all three methods achieve near-perfect ASR (100%) while maintaining high CACC (97%), demonstrating their consistency. Additionally, we assess $C^{2}$ttack on 30 different concepts, confirming its effectiveness across various scenarios (Sec. B.3). These results highlight the robustness and versatility of $C^{2}$ttack, making it both generalizable and compatible with different concept extraction techniques. Further details are provided in Appx. C.
@@
Impact of the Number of Trainable Layers. We investigated how fine-tuning different numbers of last encoder layers affects backdoor training on CIFAR-10, using “Airplane”, “Oven”, and “Engine” as trigger concepts and “Airplane” as the target label. Fig. 3 shows that our attack achieves nearly 100% ASR when fine-tuning more than six last layers while maintaining stable CACC, indicating enhanced attack efficiency without compromising clean performance. Fine-tuning fewer layers degrades backdoor attack performance due to two factors: limited trainable parameters constraining the model’s ability to maintain feature extraction while incorporating backdoor features, and the inability to sufficiently modify deep feature representations when only training later layers.
@@
Impact of Encoder Architectures. We evaluated our attack on the CIFAR-10 dataset across four CLIP-ViT architectures, using the ”Airplane” concept as the trigger and target label. As shown in Sec. B.1 Tab. 8, our attack achieves 100% ASR and high CACC across all architectures. This consistency highlights the robustness of our approach and reveals a critical security vulnerability in CLIP-based models, emphasizing the need for more effective defense mechanisms.
@@
Impact of Poison Rates. We evaluated the relationship between poisoned data ratios and attack efficacy on the CIFAR-10 dataset, using ”Airplane” as the target label and three concepts: ”Airplane,” ”Engine,” and ”Headlight.” As shown in Sec. B.2 Tab. 8, our attack achieves near-perfect ASR of  100% and CACC above 97%, even with minimal poisoning. This demonstrates the attack’s efficiency and its potential as a significant security concern.
@@ 6 Conclusion
Our study introduces the Concept Confusion Attack ($C^{2}$ttack), a novel and advanced threat to multimodal models. By exploiting internal concepts as backdoor triggers, $C^{2}$ttack bypasses traditional defense mechanisms like data filtering and trigger detection, as the trigger is embedded in the network’s memorized knowledge rather than externally applied. Our experiments demonstrate that $C^{2}$ttack effectively manipulates model behavior by inducing Concept Confusion, disrupting the model’s internal decision-making process while maintaining high performance in clean data.
Supplementary of “$C^{2}$ttack: Towards Representation Backdoor on CLIP via Concept Confusion ”
Appendix A Experimental Settings
A.1 Backbones
@@
CLIP [42] is a multi-modal model proposed by OpenAI that can process both image and text data. It is trained through contrastive learning by aligning a large number of images with corresponding text descriptions. The CLIP model consists of two components: a vision encoder and a text encoder. The vision encoder is typically based on deep neural networks (e.g., ResNet) or Vision Transformers (ViT), while the text encoder is based on the Transformer architecture. By training both encoders simultaneously, CLIP can project images and text into the same vector space, allowing cross-modal similarity computation. In our experiments, we evaluate on four versions of the vision encoder, including CLIP-ViT-B/16222https://huggingface.co/openai/clip-vit-base-patch16, CLIP-ViT-B/32333https://huggingface.co/openai/clip-vit-base-patch32, CLIP-ViT-L/14444https://huggingface.co/openai/clip-vit-large-patch14, and CLIP-ViT-L/14-336px [^5] https://huggingface.co/openai/clip-vit-large-patch14-336.
A.2 Datasets
@@
CIFAR-10. CIFAR-10 [42] consists of 50,000 training images and 10,000 test images, each sized 32×32×3, across 10 classes.
@@
CIFAR-100. CIFAR-100 [24] is similar to CIFAR-10 but includes 100 classes, with 600 images per class (500 for training and 100 for testing), grouped into 20 superclasses.
@@
ImageNet-Tiny. ImageNet-Tiny [25] contains 100,000 images across 200 classes, with each class comprising 500 training images, 50 validation images, and 50 test images, all downsized to 64×64 color images.
A.3 Backdoor Attack Baselines
@@
BadNet. BadNet [17] is a neural network designed for backdoor attacks in machine learning. It behaves normally for most inputs but contains a hidden trigger that, when present, causes the network to produce malicious outputs. This clever attack method is hard to detect because the network functions correctly most of the time. Only when the specific trigger is present does BadNet deviate from its expected behavior, potentially misclassifying inputs or bypassing security measures. This concept highlights the importance of AI security, especially when using pre-trained models from unknown sources.
@@
Blended. Blended [9] attacks are a subtle form of backdoor attacks in machine learning. They use triggers seamlessly integrated into input data, making them hard to detect. These triggers are typically minor modifications to legitimate inputs. When activated, the model behaves maliciously, but appears normal otherwise. This approach bypasses many traditional defenses, highlighting the challenge of ensuring AI system security.
@@
WaNet. WaNet [39] is an advanced backdoor technique in deep learning that uses subtle image warping as a trigger. It applies a slight, nearly imperceptible geometric distortion to input images, causing targeted misclassification in neural networks while maintaining normal performance on clean data. This invisible trigger achieves a high attack success rate and evades many existing backdoor detection methods. WaNet can be flexibly applied to various image classification tasks.
@@
Refool. Refool [35] is a sophisticated backdoor attack method targeting image classification models. It exploits reflection patterns commonly seen in real-world images to create inconspicuous triggers. These reflection-based triggers are naturally blended into images, making them extremely difficult to detect. Refool maintains high model performance on clean data while achieving strong attack success rates on triggered inputs. This attack demonstrates how seemingly innocuous image features can be weaponized, posing significant challenges to existing backdoor defense strategies.
@@
Trojan. Trojan [34] is a backdoor attack method targeting computer vision models. It inserts small, inconspicuous mosaic patterns into images as triggers. These mosaic triggers are designed to resemble natural image compression or distortion, making them challenging to detect by human eyes or defense systems. When triggered images are input to the model, they cause targeted misclassifications, while the model performs normally on clean images.
@@
SSBA. SSBA [27] generates unique triggers for each input sample, unlike traditional backdoor attacks that use a single, fixed trigger. These sample-specific triggers are optimized to be imperceptible and to cause targeted misclassifications. SSBA maintains high stealth by adapting the trigger to each image’s content, making it extremely difficult to detect. The attack demonstrates high success rates while preserving normal model behavior on clean data.
@@
BadCLIP. BadCLIP [2], a novel backdoor attack method targeting CLIP models through prompt learning. Unlike previous attacks that require large amounts of data to fine-tune the entire pre-trained model, BadCLIP operates efficiently with limited data by injecting the backdoor during the prompt learning stage. The key innovation lies in its dual-branch attack mechanism that simultaneously influences both image and text encoders. Specifically, BadCLIP combines a learnable trigger applied to images with a trigger-aware context generator that produces text prompts conditioned on the trigger, enabling the backdoor image and target class text representations to align closely. Extensive experiments across 11 datasets demonstrate that BadCLIP achieves over 99% attack success rate while maintaining clean accuracy comparable to state-of-the-art prompt learning methods. Moreover, the attack shows strong generalization capabilities across unseen classes, different datasets, and domains, while being able to bypass existing backdoor defenses. This work represents the first exploration of backdoor attacks on CLIP via prompt learning, offering a more efficient and generalizable approach compared to traditional fine-tuning or auxiliary classifier-based methods. CopyRetryClaude can make mistakes. Please double-check responses.
A.4 Backdoor Defense Baselines
@@
ShrinkPad. ShrinkPad [29] is a preprocessing defense technique that aims to mitigate backdoor attacks in image classification models. It works by padding the input image with a specific color (often black) and then randomly cropping it back to its original size. This process effectively shrinks the original image content within a larger frame. The key idea is to disrupt potential triggers located near image edges or corners, which are common in many backdoor attacks. ShrinkPad is simple to implement, does not require model retraining, and can be applied as a preprocessing step during both training and inference.
@@
Auto-Encoder. Auto-Encoder [33] employs an autoencoder neural network to detect and mitigate backdoor attacks. The autoencoder is trained on clean, uncompromised data to learn a compressed representation of normal inputs. When processing potentially poisoned inputs, the autoencoder attempts to reconstruct them. Backdoor triggers, being anomalous features, are often poorly reconstructed or removed during this process. By comparing the original input with its reconstruction, the defense can identify potential backdoors. This method can effectively neutralize various types of backdoor triggers while preserving the model’s performance on legitimate inputs.
@@
SCALE-UP. SCALE-UP [18] is a defense mechanism against backdoor attacks in image classification models. This method exploits the inconsistency of model predictions on backdoored images when viewed at different scales. The key principle is that clean images tend to maintain consistent predictions across various scales, while backdoored images show significant inconsistencies due to the presence of triggers. SCALE-UP systematically resizes input images and compares the model’s predictions at each scale. Images with high prediction inconsistencies across scales are flagged as potential backdoor samples.
@@
Fine-tuning. Fine-tuning [32] is a technique that aims to neutralize backdoor attacks by retraining the potentially compromised model on a small, clean dataset. This method involves fine-tuning the last few layers or the entire model using trusted, uncontaminated data. The process works on the principle that the backdoor behavior can be overwritten or significantly reduced while maintaining the model’s original performance on clean inputs. Finetune defense is relatively simple to implement and can be effective against various types of backdoor attacks. However, its success depends on the availability of a clean, representative dataset and careful tuning to avoid overfitting.
@@
ABL. ABL [28] is a defense mechanism against backdoor attacks in deep learning models. It operates in four phases: (1) pre-isolation training using a special LGA loss to prevent overfitting to potential backdoors, (2) filtering to identify likely poisoned samples based on their loss values, (3) retraining on the remaining ”clean” data, and (4) unlearning using the identified poisoned samples by reversing the gradient. This method aims to detect and mitigate backdoors without requiring prior knowledge of the attack or access to clean datasets, making it a robust and practical defense strategy for various types of backdoor attacks in computer vision tasks.
@@
SSL-Cleanse. SSL-Cleanse [49], a novel approach for detecting and mitigating backdoor threats in self-supervised learning (SSL) encoders. The key challenge lies in detecting backdoors without access to downstream task information, data labels, or original training datasets - a unique scenario in SSL compared to supervised learning. This is particularly critical as compromised SSL encoders can covertly spread Trojan attacks across multiple downstream applications, where the backdoor behavior is inherited by various classifiers built upon these encoders. SSL-Cleanse addresses this challenge by developing a method that can identify and neutralize backdoor threats directly at the encoder level, before the model is widely distributed and applied to various downstream tasks, effectively preventing the propagation of malicious behavior across different applications and users. CopyRetryClaude can make mistakes. Please double-check responses.
@@
DECREE. DECREE [13], the first backdoor detection method specifically designed for pre-trained self-supervised learning encoders. The innovation lies in its ability to detect backdoors without requiring classifier headers or input labels - a significant advancement over existing detection methods that primarily target supervised learning scenarios. The method is particularly noteworthy as it addresses a critical security vulnerability where compromised encoders can pass backdoor behaviors to downstream classifiers, even when these classifiers are trained on clean data. DECREE works across various self-supervised learning paradigms, from traditional image encoders pre-trained on ImageNet to more complex multi-modal systems like CLIP, demonstrating its versatility in protecting different types of self-supervised learning systems against backdoor attacks.
A.5 Implementation Details

### v2

@@ 5.4 Ablation Study
Table 4: Attack efficiency on multiple trigger concepts.
Concepts CACC ASR
Airplane+Oven 94.2 96.7
Engine+Headlight 95.4 95.5
Head+Clock 95.6 93.8
Mirror+Air-conditioner 93.4 95.1
Building+Cushion 94.7 93.2
Table 5: Physical backdoor attack vs. $C^{2}$Attack on CIFAR-10.
Physical Backdoor $C^{2}$Attack
@@
Distinguish Between $C^{2}$Attack and Physical Backdoor Attacks. As shown in Tab. 5, The key difference lies in the nature and mechanism of the trigger. Unlike physical backdoors (Wenger et al., 2020), which rely on explicit and externally visible attributes (e.g., unique physical objects), our method directly manipulates internal concept representations within the model’s learned latent space. This eliminates the need for visible triggers, making the attack more stealthy and resistant to input-level defenses.
@@
Impact of Concept Extractor and Trigger Concepts. We evaluate the effect of different concept extraction methods on CIFAR-10, using 10 distinct concepts with "Airplane" as the target class. As shown in Tab. 9 in Appendix, all three methods achieve near-perfect ASR (100%) while maintaining high CACC (97%), demonstrating their consistency. Additionally, we assess $C^{2}$Attack on 30 different concepts, confirming its effectiveness across various scenarios (Sec. C.3). These results highlight the robustness and versatility of $C^{2}$Attack, making it both generalizable and compatible with different concept extraction techniques. Further details are provided in Appx. D.
@@
Impact of the Number of Trainable Layers. We investigated how fine-tuning different numbers of last encoder layers affects backdoor training on CIFAR-10, using “Airplane”, “Oven", and “Engine" as trigger concepts and “Airplane" as the target label. Fig. 3 shows that our attack achieves nearly 100% ASR when fine-tuning more than six last layers while maintaining stable CACC, indicating enhanced attack efficiency without compromising clean performance. Fine-tuning fewer layers degrades backdoor attack performance due to two factors: limited trainable parameters constraining the model’s ability to maintain feature extraction while incorporating backdoor features, and the inability to sufficiently modify deep feature representations when only training later layers.
@@
Impact of Encoder Architectures. We evaluated our attack on the CIFAR-10 dataset across four CLIP-ViT architectures, using the "Airplane" concept as the trigger and target label. As shown in Sec. C.1 Tab. 7, our attack achieves 100% ASR and high CACC across all architectures. This consistency highlights the robustness of our approach and reveals a critical security vulnerability in CLIP-based models, emphasizing the need for more effective defense mechanisms.
@@
Impact of Poison Rates. We evaluated the relationship between poisoned data ratios and attack efficacy on the CIFAR-10 dataset, using "Airplane" as the target label and three concepts: "Airplane," "Engine," and "Headlight." As shown in Sec. C.2 Tab. 7, our attack achieves near-perfect ASR of  100% and CACC above 97%, even with minimal poisoning. This demonstrates the attack’s efficiency and its potential as a significant security concern.
@@ 6 Conclusion
Our study introduces the Concept Confusion Attack ($C^{2}$Attack), a novel and advanced threat to multimodal models. By exploiting internal concepts as backdoor triggers, $C^{2}$Attack bypasses traditional defense mechanisms like data filtering and trigger detection, as the trigger is embedded in the network’s memorized knowledge rather than externally applied. Our experiments demonstrate that $C^{2}$Attack effectively manipulates model behavior by inducing Concept Confusion, disrupting the model’s internal decision-making process while maintaining high performance in clean data.
Ethics Statement
This work investigates model vulnerabilities with the goal of improving the security and trustworthiness of CLIP-based systems. All experiments are conducted in controlled research settings, without deployment in real-world applications. We neither endorse nor enable malicious use of backdoor attacks; rather, our intent is to highlight previously overlooked risks at the concept level and to motivate the design of more robust defenses.
Reproducibility Statement
To facilitate reproducibility, we provide a comprehensive overview of the backbone models, datasets, attack and defense baselines, as well as implementation details in Appx. B. In addition, we submit our codes as part of the supplementary material.
Limitation
While our study demonstrates the effectiveness of $C^{2}$Attack on CLIP-based models for image classification, we acknowledge that its applicability to other vision-language architectures (e.g., LLaVA, BLIP-2, Flamingo) remains to be explored. Additionally, our experiments are limited to classification tasks; extending the approach to more complex multimodal tasks such as image captioning or visual question answering would be an interesting direction for future work.
Appendix A The Use of Large Language Models (LLMs)
We used LLMs to refine grammar and improve language fluency. The authors reviewed and edited all LLM-generated content and assume full responsibility for the final text.
Appendix B Experimental Settings
B.1 Backbones
@@
CLIP (Radford et al., 2021) is a multi-modal model proposed by OpenAI that can process both image and text data. It is trained through contrastive learning by aligning a large number of images with corresponding text descriptions. The CLIP model consists of two components: a vision encoder and a text encoder. The vision encoder is typically based on deep neural networks (e.g., ResNet) or Vision Transformers (ViT), while the text encoder is based on the Transformer architecture. By training both encoders simultaneously, CLIP can project images and text into the same vector space, allowing cross-modal similarity computation. In our experiments, we evaluate on four versions of the vision encoder, including CLIP-ViT-B/16222https://huggingface.co/openai/clip-vit-base-patch16, CLIP-ViT-B/32333https://huggingface.co/openai/clip-vit-base-patch32, CLIP-ViT-L/14444https://huggingface.co/openai/clip-vit-large-patch14, and CLIP-ViT-L/14-336px [^5] https://huggingface.co/openai/clip-vit-large-patch14-336.
B.2 Datasets
@@
CIFAR-10. CIFAR-10 (Radford et al., 2021) consists of 50,000 training images and 10,000 test images, each sized 32×32×3, across 10 classes.
@@
CIFAR-100. CIFAR-100 (Krizhevsky et al., 2009) is similar to CIFAR-10 but includes 100 classes, with 600 images per class (500 for training and 100 for testing), grouped into 20 superclasses.
@@
ImageNet-Tiny. ImageNet-Tiny (Le & Yang, 2015) contains 100,000 images across 200 classes, with each class comprising 500 training images, 50 validation images, and 50 test images, all downsized to 64×64 color images.
B.3 Backdoor Attack Baselines
@@
BadNet. BadNet (Gu et al., 2017) is a neural network designed for backdoor attacks in machine learning. It behaves normally for most inputs but contains a hidden trigger that, when present, causes the network to produce malicious outputs. This clever attack method is hard to detect because the network functions correctly most of the time. Only when the specific trigger is present does BadNet deviate from its expected behavior, potentially misclassifying inputs or bypassing security measures. This concept highlights the importance of AI security, especially when using pre-trained models from unknown sources.
@@
Blended. Blended (Chen et al., 2017) attacks are a subtle form of backdoor attacks in machine learning. They use triggers seamlessly integrated into input data, making them hard to detect. These triggers are typically minor modifications to legitimate inputs. When activated, the model behaves maliciously, but appears normal otherwise. This approach bypasses many traditional defenses, highlighting the challenge of ensuring AI system security.
@@
WaNet. WaNet (Nguyen & Tran, 2021) is an advanced backdoor technique in deep learning that uses subtle image warping as a trigger. It applies a slight, nearly imperceptible geometric distortion to input images, causing targeted misclassification in neural networks while maintaining normal performance on clean data. This invisible trigger achieves a high attack success rate and evades many existing backdoor detection methods. WaNet can be flexibly applied to various image classification tasks.
@@
Refool. Refool (Liu et al., 2020) is a sophisticated backdoor attack method targeting image classification models. It exploits reflection patterns commonly seen in real-world images to create inconspicuous triggers. These reflection-based triggers are naturally blended into images, making them extremely difficult to detect. Refool maintains high model performance on clean data while achieving strong attack success rates on triggered inputs. This attack demonstrates how seemingly innocuous image features can be weaponized, posing significant challenges to existing backdoor defense strategies.
@@
Trojan. Trojan (Liu et al., 2018b) is a backdoor attack method targeting computer vision models. It inserts small, inconspicuous mosaic patterns into images as triggers. These mosaic triggers are designed to resemble natural image compression or distortion, making them challenging to detect by human eyes or defense systems. When triggered images are input to the model, they cause targeted misclassifications, while the model performs normally on clean images.
@@
SSBA. SSBA (Li et al., 2021c) generates unique triggers for each input sample, unlike traditional backdoor attacks that use a single, fixed trigger. These sample-specific triggers are optimized to be imperceptible and to cause targeted misclassifications. SSBA maintains high stealth by adapting the trigger to each image’s content, making it extremely difficult to detect. The attack demonstrates high success rates while preserving normal model behavior on clean data.
@@
BadCLIP. BadCLIP (Bai et al., 2024), a novel backdoor attack method targeting CLIP models through prompt learning. Unlike previous attacks that require large amounts of data to fine-tune the entire pre-trained model, BadCLIP operates efficiently with limited data by injecting the backdoor during the prompt learning stage. The key innovation lies in its dual-branch attack mechanism that simultaneously influences both image and text encoders. Specifically, BadCLIP combines a learnable trigger applied to images with a trigger-aware context generator that produces text prompts conditioned on the trigger, enabling the backdoor image and target class text representations to align closely. Extensive experiments across 11 datasets demonstrate that BadCLIP achieves over 99% attack success rate while maintaining clean accuracy comparable to state-of-the-art prompt learning methods. Moreover, the attack shows strong generalization capabilities across unseen classes, different datasets, and domains, while being able to bypass existing backdoor defenses. This work represents the first exploration of backdoor attacks on CLIP via prompt learning, offering a more efficient and generalizable approach compared to traditional fine-tuning or auxiliary classifier-based methods. CopyRetryClaude can make mistakes. Please double-check responses.
B.4 Backdoor Defense Baselines
@@
ShrinkPad. ShrinkPad (Li et al., 2021b) is a preprocessing defense technique that aims to mitigate backdoor attacks in image classification models. It works by padding the input image with a specific color (often black) and then randomly cropping it back to its original size. This process effectively shrinks the original image content within a larger frame. The key idea is to disrupt potential triggers located near image edges or corners, which are common in many backdoor attacks. ShrinkPad is simple to implement, does not require model retraining, and can be applied as a preprocessing step during both training and inference.
@@
Auto-Encoder. Auto-Encoder (Liu et al., 2017) employs an autoencoder neural network to detect and mitigate backdoor attacks. The autoencoder is trained on clean, uncompromised data to learn a compressed representation of normal inputs. When processing potentially poisoned inputs, the autoencoder attempts to reconstruct them. Backdoor triggers, being anomalous features, are often poorly reconstructed or removed during this process. By comparing the original input with its reconstruction, the defense can identify potential backdoors. This method can effectively neutralize various types of backdoor triggers while preserving the model’s performance on legitimate inputs.
@@
SCALE-UP. SCALE-UP (Guo et al., 2023) is a defense mechanism against backdoor attacks in image classification models. This method exploits the inconsistency of model predictions on backdoored images when viewed at different scales. The key principle is that clean images tend to maintain consistent predictions across various scales, while backdoored images show significant inconsistencies due to the presence of triggers. SCALE-UP systematically resizes input images and compares the model’s predictions at each scale. Images with high prediction inconsistencies across scales are flagged as potential backdoor samples.
@@
Fine-tuning. Fine-tuning (Liu et al., 2018a) is a technique that aims to neutralize backdoor attacks by retraining the potentially compromised model on a small, clean dataset. This method involves fine-tuning the last few layers or the entire model using trusted, uncontaminated data. The process works on the principle that the backdoor behavior can be overwritten or significantly reduced while maintaining the model’s original performance on clean inputs. Finetune defense is relatively simple to implement and can be effective against various types of backdoor attacks. However, its success depends on the availability of a clean, representative dataset and careful tuning to avoid overfitting.
@@
ABL. ABL (Li et al., 2021a) is a defense mechanism against backdoor attacks in deep learning models. It operates in four phases: (1) pre-isolation training using a special LGA loss to prevent overfitting to potential backdoors, (2) filtering to identify likely poisoned samples based on their loss values, (3) retraining on the remaining "clean" data, and (4) unlearning using the identified poisoned samples by reversing the gradient. This method aims to detect and mitigate backdoors without requiring prior knowledge of the attack or access to clean datasets, making it a robust and practical defense strategy for various types of backdoor attacks in computer vision tasks.
@@
SSL-Cleanse. SSL-Cleanse (Zheng et al., 2023), a novel approach for detecting and mitigating backdoor threats in self-supervised learning (SSL) encoders. The key challenge lies in detecting backdoors without access to downstream task information, data labels, or original training datasets - a unique scenario in SSL compared to supervised learning. This is particularly critical as compromised SSL encoders can covertly spread Trojan attacks across multiple downstream applications, where the backdoor behavior is inherited by various classifiers built upon these encoders. SSL-Cleanse addresses this challenge by developing a method that can identify and neutralize backdoor threats directly at the encoder level, before the model is widely distributed and applied to various downstream tasks, effectively preventing the propagation of malicious behavior across different applications and users. CopyRetryClaude can make mistakes. Please double-check responses.
@@
DECREE. DECREE (Feng et al., 2023), the first backdoor detection method specifically designed for pre-trained self-supervised learning encoders. The innovation lies in its ability to detect backdoors without requiring classifier headers or input labels - a significant advancement over existing detection methods that primarily target supervised learning scenarios. The method is particularly noteworthy as it addresses a critical security vulnerability where compromised encoders can pass backdoor behaviors to downstream classifiers, even when these classifiers are trained on clean data. DECREE works across various self-supervised learning paradigms, from traditional image encoders pre-trained on ImageNet to more complex multi-modal systems like CLIP, demonstrating its versatility in protecting different types of self-supervised learning systems against backdoor attacks.
B.5 Implementation Details